A car travelling at 20 m/s collides with a stationary lorry and is brought to rest in just 0.02s, Calculate the deceleration of the car.


1
Expert's answer
2021-01-20T07:23:44-0500

We can find the deceleration of the car from the formula:


"a=\dfrac{v-v_0}{t},""a=\dfrac{0-20\ \dfrac{m}{s}}{0.02\ s}=-1000\ \dfrac{m}{s^2}."

Answer:

"a=-1000\ \dfrac{m}{s^2}."